# SCAIL: Towards Studio-Grade Character Animation via In-Context Learning of 3D-Consistent Pose Representations

### v1

@@ Abstract
Achieving character animation that meets studio-grade production standards remains challenging despite recent progress. Existing approaches can transfer motion from a driving video to a reference image, but often fail to preserve structural fidelity and temporal consistency in wild scenarios involving complex motion and cross-identity animations. In this work, we present SCAIL (Studio-grade Character Animation via In-context Learning), a framework designed to address these challenges from two key innovations. First, we propose a novel 3D pose representation, providing a more robust and flexible motion signal. Second, we introduce a full-context pose injection mechanism within a diffusion-transformer architecture, enabling effective spatio-temporal reasoning over full motion sequences. To align with studio-level requirements, we develop a curated data pipeline ensuring both diversity and quality, and establish a comprehensive benchmark for systematic evaluation. Experiments show that SCAIL achieves state-of-the-art performance and advances character animation toward studio-grade reliability and realism. Code and model will be publicly available at our project page.
@@ 1 Introduction
High-fidelity character animation has tremendous potential for film production. Conventional filmmaking pipelines rely on a complicated workflow—involving motion capture, rigging, rendering—that demands expensive hardware and significant expert labor. Recently, the emergence of video generation models [VDM, SVD, cogvideox, Wan] introduces a new paradigm: given a reference image and a driving video, such models can directly synthesize a photorealistic animation that follows the target motion in driving video, dramatically lowering the production barrier for high-quality character animation. Similar to real-world production pipelines, current methods [MagicAnimate, AnimateAnyone, champ, MimicMotion, Stableanimator, animate-x, Unianimate-dit, Wan-Animate] typically begin by extracting skeletal motion sequences from the driving video as a form of “motion capture”, and then inject this information into a video generation model to perform “rigging” and “rendering”. However, in studio-grade production setting, these methods often struggle with challenging scenarios, such as complex motions (e.g., turning, rolling, flipping), multi-person interactions (e.g., dancing, hugging, fighting), and cross-domain animation where the reference and driving subjects differ significantly in appearance or body shape. As a result, generated videos frequently exhibit artifacts including distorted appearance, implausible body poses, and incorrect occlusions. These limitations reveal that current skeletal pose representations and generation models fail to adequately capture 3D structure of driving motions, inter-character spatial and occlusion relationships, and temporal correlations of motion sequences.
@@
To address the above limitations, we present SCAIL, a framework that revisits two core technical bottlenecks: (1) how to build a pose representation that effectively bridges driving and generated videos with unambiguous and accurate motion encoding, and (2) how to inject pose control in a way that enables the model to capture spatio-temporal motion structures. First, for motion representation, prior works [AnimateAnyone, MimicMotion, Stableanimator, Unianimate-dit, VACE, Wan-Animate] typically rely on 2D skeletons (extracted from DWPose [DWPose], ViTPose [ViTPose], etc.), which suffer from noisy predictions in complex motions and cannot encode occlusion. Some recent methods [champ, Vividpose] adopt SMPL-based [SMPL] controls, which offer strong 3D human priors but have limited augmentation flexibility and cause identity leakage due to person-specific shape parameters. To address these issues, we propose a novel 3D pose representation. Specifically, we connect estimated 3D human keypoints according to skeletal topology and represent bones as spatial cylinders. The resulting 3D skeleton is rasterized to obtain 2D motion guidance signals. Building on this representation, we further design pose augmentation and retargeting strategies, enabling seamless scaling across diverse characters and scenarios. Second, regarding pose injection, a common approach in diffusion transformer (DiT) [DiT] model is to concatenate conditions channel-wise [realis-dance, Wan-Animate]. However, this only provides local motion cues without capturing global motion dependencies. To overcome this, we propose a full-context pose injection mechanism within a DiT-based architecture. This design allows the model to attend to the entire pose sequence when generating each frame, enabling it to reason about motion context across time, and better capture high-level motion semantics. As a result, our method supports robust motion transfer across diverse body types and visual domains.
@@ 2.1 Diffusion Models for Video Generation
Diffusion models [ddpm, ddim] effectively overcome the training instability and mode collapse of generative adversarial networks (GANs) [gan], and emerge as the dominant paradigm in visual content generation. The success of the Stable Diffusion (SD) [SD, SDXL] in image synthesis has naturally inspired the extension from image to video. Subsequently, the Diffusion Transformer (DiT) [DiT] architecture, combined with RoPE [rope] for position encoding, offers superior modeling capability and scalability, becoming the leading backbone for high-quality video generation [sora, kling, cogvideox, Wan]. More recently, diffusion-based video generation has advanced toward controllability, enabling control over camera viewpoints [cameractrl, camco], motion trajectories [Tora, 3DTrajMaster], and scene structures [Gen1, make-your-video, DrivingDiffusion]. In this work, we focus on character video generation with precise pose control.
@@ 2.2 Character Image Animation
Character image animation aims to generate photorealistic and temporally coherent video, where the appearance remains consistent with a given reference image and the motion follows a driving video. AnimateAnyone [AnimateAnyone] extracts 2D skeletons from the driving video as motion guidance, and designs Pose Guider and ReferenceNet modules to control motion and appearance. Subsequent works explore various improvements. Champ [champ] integrates multiple motion signals, including depth maps, SMPL [SMPL] normals, and 2D skeletons. Animate-X [animate-x] introduces a skeleton augmentation strategy to enable animation of anthropomorphic characters with large body-ratio discrepancies. Moreover, MimicMotion [MimicMotion] address hand and facial distortions via regional loss. DanceTogether [DanceTogether] extends this line of research to multi-person animation. UniAnimate-DiT [Unianimate-dit], VACE [VACE] and Wan-Animate [Wan-Animate] replace U-Net with Transformer as the denoising backbone, significantly improving the generation quality. Despite progress, current models still struggle with complex scenarios and cross-pair animations, which hinders their deployment in real production.
@@ 3.1 Preliminaries
Latent Diffusion Models. Latent diffusion models [SD, latent-shift] reduce the computational cost of pixel-space diffusion [ddpm, Imagen] by operating in a compressed latent space. Given an input video $\bm{x}$, a pretrained VAE encoder [Wan] $\mathcal{E}$ first maps it into a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$. During the forward diffusion process, Gaussian noise is progressively added over $T$ timesteps, formulated as:
@@
Diffusion Transformer (DiT). Traditional diffusion models typically rely on U-Net as denoising backbones. Recent works propose Diffusion Transformer (DiT) [DiT] architecture, which supports variable input resolution and sequence length. The DiT combines the generative power of diffusion models with the flexibility of Transformers [transformer], enabling better scalability and stronger modeling capacity.
@@
3D-Consistent Pose Representation. Recent methods [AnimateAnyone, animate-x, Unianimate-dit, VACE, Wan-Animate] that employ 2D keypoint-based skeletons as motion controls perform well on simple actions. However, they often struggle under complex studio-level production scenarios, leading to abnormal human kinematics and violation of motion physics. 2D keypoints cannot encode 3D occlusion, making it difficult for models to generate realistic motion videos from incomplete or ambiguous pose signals. Another challenge arises from the discrepancy between the reference image and driving pose, which often differ in body shape or identity. Previous attempts [animate-x, Unianimate-dit] mitigate the discrepancy through 2D skeleton scaling in training or heuristic retargeting in testing, yet 2D-derived adaptation inherently suffers from deformation due to inconsistency with 3D motion dynamics and further amplifies estimation noise and inaccuracies.
We argue that an ideal driving pose representation should perceive depth and occlusion, remain identity-agnostic, and retain the potential to scale to multi-person settings. To achieve that, we employ NLFPose [nlfpose] to estimate 3D body keypoints $\mathcal{J}=\{\text{J}_{i,j}\}_{i=1,j=1}^{T,N}$, where $\tilde{\text{J}}_{i,j}$ denotes the predicted 3D joint $j$ at frame $i$, and connect them according to the skeletal topology in 3D space. We discard SMPL mesh-like representations, which are inherently human-specific and offer limited controllability for augmentation or retargeting, often resulting in identity leakage. To retain occlusion and depth cues lost in 2D projection, we render 3D skeletons as cylindrical segments to provide motion-driving signals.
@@
Full-Context Driving Pose Injection. Our model builds upon a DiT-based Image-to-Video (I2V) model [Wan] and inject pose control signals for motion guidance. The common practice for pose information injection is through channel concat [realis-dance, Wan-Animate] or using a pose adapter [Unianimate-dit], where the driving pose sequence is embedded and added to the noisy video latents. We first implement the channel concat approach as illustrated in Figure 4, following [realis-dance]. While this method shows decent pose-following capability in generation, we find it tend to generate unnatural human pose when the motion is complex despite conditioned on accurate and unambiguous pose. A typical example is the turning motion, where the model often fails to correctly distinguish between front and back views, generating awkward postures. We assume that this limitation stems from the per-frame addition scheme, which fails to provide sufficient temporal context for motion understanding. Since motion is inherently sequential, many actions can only be interpreted correctly within a temporal context. Therefore, we design a new strategy termed full-context pose injection, which enables the model to reason over the entire pose sequence. Specifically, we directly concatenates conditional pose tokens with noisy video tokens to facilitate spatial-temporal interactions between the two modality. As shown in Figure 4, this strategy handles motion correctly in complex scenarios.
@@
where $\text{shift}_{W}$ is a constant shift magnitude and $W_{\max}$ is the width of the reference image. As show in Figure 3, this design helps the model to distinguish driving pose tokens from reference image tokens and noisy video tokens in conjunction with the modified I2V Mask and reference image token injection strategy, enhancing the overall performance. Considering the modality gap between the noisy video tokens and the driving pose tokens, we observe that the model performs best particularly when the shift distance is a relatively large constant. To accommodate different downsampling ratios for pose context, mean-pooling on the 3D-RoPE [rope] frequencies is performed according to the applied ratio.
@@ 3.3 Data Collection
Filtering and Pose Extraction Pipeline. We develop a curated pipeline to construct a high-quality dataset for character animation. To meet the requirements for generating studio-grade character animation, we have established the following criteria for data filtering: (1) the character should be the primary focus of each frame, (2) the person should exhibit explicit motion, and (3) the pose should be visually complete, covering either the upper body or the full body. We first detect human presence using YOLO [yolo] and discard clips where characters are not the main subjects. Next, 2D keypoints are extracted using DWPose [DWPose], retaining only half-body or full-body videos. After this stage, multi-person videos can be further separated based on the number of bounding boxes detected by YOLO [yolo]. We further employ a large language model [glm] to analyze captions, complementing the complex rule-based logic of multi-person motion bounding box checking. For multi-person videos, we employ Samurai [samurai] to track and split the mask of each character, generating multiple single-human video splits. Subsequently, 3D keypoints are extracted using NLFPose [nlfpose] for each video that contain one main character. To meet the motion criteria, we compute the motion speed from the estimated 3D keypoints and discard samples with limited motion, resulting in a clean and motion-rich training set. The motion speed $v$ is calculated as:
@@
Finally, The 3D skeletons are represented as cylindrical segments in 3D space and rendered using the Taichi [taichi] renderer to preserve spatial relationships. For multi-person data, each split is independently processed by NLFPose to estimate 3D skeletons, which are then composed together in 3D space and rasterized on the 2D canvas. This procedure preserves inter-person occlusion relationships, benefiting from NLFPose’s accurate depth estimation. In practice, we find that our proposed multi-stage pipeline provides more accurate estimation than direct multi-person pose estimation methods like PromptHMR [PromptHMR], especially in wild cases involving complex interaction.
@@
UniAnimate-DiT [Unianimate-dit] 17.79 0.637 0.242 362.27 2.5% 1.7% 0.8% 1.7%
VACE [VACE] 16.73 0.588 0.263 264.71 9.2% 14.2% 18.3% 32.5%
Wan-Animate [Wan-Animate] 18.54 0.648 0.221 187.61 35.0% 28.3% 24.2% 20.0%
@@
Table 1: Quantitative comparison for SCAIL-14B and baselines. All compared methods are built upon 14B Wan [Wan] foundation models.
@@ 3.4 Studio-Bench for Evaluation
To evaluate the model capabilities under studio-grade standards, we propose Studio-Bench, a new evaluation benchmark tailored for Studio-Grade character animation. Previous evaluations [MagicAnimate, AnimateAnyone, champ, MimicMotion, Stableanimator, Unianimate-dit] primarily focus on simple actions, which fail to capture the challenges present in film production, such as complex and dynamic motions, multi-person interactions, and cross-domain animations. Specifically, our benchmark consists of two parts: Self-Driven and Cross-Driven. The first part evaluates motion adherence under complex actions. We examine whether the generated videos maintain correct body structure during large-scale motions, and whether the model can accurately capture inter-person relationships in multi-character scenes, totally containing 130 clips. Quantitative metrics can be computed by directly comparing generated results with paired ground-truth videos for this subset. The second part measures the model’s transferability when the reference image and driving video differ in identity or domain. We manually select a set of real reference images to test generalization, and further construct additional reference images using an image generation tool [seedream] to introduce variations in style and body shape, obtaining 120 single-character pairs and 10 multi-character pairs. As we simulate in-the-wild reference-driving discrepancy, pose retarget should be enabled for single-character pairs in this subset. We incorporate the retarget logic of [Unianimate-dit] for VACE [VACE] which do not natively support retargeting. For this subset, we conduct user study to assess the generation quality. All quantitative results on the subset are based on the 120 single-character pairs to ensure fairness considering some baselines’ incompatibility with multi-character settings. Evaluation samples are strictly excluded from the training data.

### v2

@@ Abstract
Achieving character animation that meets studio-grade production standards remains challenging despite recent progress. Existing approaches can transfer motion from a driving video to a reference image, but often fail to preserve structural fidelity and temporal consistency in wild scenarios involving complex motion and cross-identity animations. In this work, we present SCAIL (a framework toward Studio-grade Character Animation via In-context Learning), a framework designed to address these challenges from two key innovations. First, we propose a novel 3D pose representation, providing a more robust and flexible motion signal. Second, we introduce a full-context pose injection mechanism within a diffusion-transformer architecture, enabling effective spatio-temporal reasoning over full motion sequences. To align with studio-level requirements, we develop a curated data pipeline ensuring both diversity and quality, and establish a comprehensive benchmark for systematic evaluation. Experiments show that SCAIL achieves state-of-the-art performance and advances character animation toward studio-grade reliability and realism. Code and model are available at zai-org/SCAIL.
@@ 1 Introduction
High-fidelity character animation has tremendous potential for film production. Conventional filmmaking pipelines rely on a complicated workflow—involving motion capture, rigging, rendering—that demands expensive hardware and significant expert labor. Recently, the emergence of video generation models [12, 2, 47, 36] introduces a new paradigm: given a reference image and a driving video, such models can directly synthesize a photorealistic animation that follows the target motion in driving video, dramatically lowering the production barrier for high-quality character animation. Similar to real-world production pipelines, current methods [44, 15, 52, 49, 31, 30, 38, 4] typically begin by extracting skeletal motion sequences from the driving video as a form of “motion capture”, and then inject this information into a video generation model to perform “rigging” and “rendering”. However, in studio-grade production setting, these methods often struggle with challenging scenarios, such as complex motions (e.g., turning, rolling, flipping), multi-person interactions (e.g., dancing, hugging, fighting), and cross-domain animation where the reference and driving subjects differ significantly in appearance or body shape. As a result, generated videos frequently exhibit artifacts including distorted appearance, implausible body poses, and incorrect occlusions. These limitations reveal that current skeletal pose representations and generation models fail to adequately capture 3D structure of driving motions, inter-character spatial and occlusion relationships, and temporal correlations of motion sequences.
@@
To address the above limitations, we present SCAIL, a framework that revisits two core technical bottlenecks: (1) how to build a pose representation that effectively bridges driving and generated videos with unambiguous and accurate motion encoding, and (2) how to inject pose control in a way that enables the model to capture spatio-temporal motion structures. First, for motion representation, prior works [15, 49, 31, 38, 17, 4] typically rely on 2D skeletons (extracted from DWPose [46], ViTPose [43], etc.), which suffer from noisy predictions in complex motions and cannot encode occlusion. Some recent methods [52, 37] adopt SMPL-based [20] controls, which offer strong 3D human priors but have limited augmentation flexibility and cause identity leakage due to person-specific shape parameters. To address these issues, we propose a novel 3D pose representation. Specifically, we connect estimated 3D human keypoints according to skeletal topology and represent bones as spatial cylinders. The resulting 3D skeleton is rasterized to obtain 2D motion guidance signals. Building on this representation, we further design pose augmentation and retargeting strategies, enabling seamless scaling across diverse characters and scenarios. Second, regarding pose injection, a common approach in diffusion transformer (DiT) [23] model is to concatenate conditions channel-wise [51, 4]. However, this only provides local motion cues without capturing global motion dependencies. To overcome this, we propose a full-context pose injection mechanism within a DiT-based architecture. This design allows the model to attend to the entire pose sequence when generating each frame, enabling it to reason about motion context across time, and better capture high-level motion semantics. As a result, our method supports robust motion transfer across diverse body types and visual domains.
@@ 2.1 Diffusion Models for Video Generation
Diffusion models [11, 28] effectively overcome the training instability and mode collapse of generative adversarial networks (GANs) [8], and emerge as the dominant paradigm in visual content generation. The success of the Stable Diffusion (SD) [25, 24] in image synthesis has naturally inspired the extension from image to video. Subsequently, the Diffusion Transformer (DiT) [23] architecture, combined with RoPE [29] for position encoding, offers superior modeling capability and scalability, becoming the leading backbone for high-quality video generation [22, 18, 47, 36]. More recently, diffusion-based video generation has advanced toward controllability, enabling control over camera viewpoints [9, 42], motion trajectories [50, 6], and scene structures [5, 41, 19]. In this work, we focus on character video generation with precise pose control.
@@ 2.2 Character Image Animation
Character image animation aims to generate photorealistic and temporally coherent video, where the appearance remains consistent with a given reference image and the motion follows a driving video. AnimateAnyone [15] extracts 2D skeletons from the driving video as motion guidance, and designs Pose Guider and ReferenceNet modules to control motion and appearance. Subsequent works explore various improvements. Champ [52] integrates multiple motion signals, including depth maps, SMPL [20] normals, and 2D skeletons. Animate-X [30] introduces a skeleton augmentation strategy to enable animation of anthropomorphic characters with large body-ratio discrepancies. Moreover, MimicMotion [49] address hand and facial distortions via regional loss. DanceTogether [3] extends this line of research to multi-person animation. UniAnimate-DiT [38], VACE [17] and Wan-Animate [4] replace U-Net with Transformer as the denoising backbone, significantly improving the generation quality. Despite progress, current models still struggle with complex scenarios and cross-pair animations, which hinders their deployment in real production.
@@ 3.1 Preliminaries
Latent Diffusion Models. Latent diffusion models [25, 1] reduce the computational cost of pixel-space diffusion [11, 10] by operating in a compressed latent space. Given an input video $\bm{x}$, a pretrained VAE encoder [36] $\mathcal{E}$ first maps it into a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$. During the forward diffusion process, Gaussian noise is progressively added over $T$ timesteps, formulated as:
@@
Diffusion Transformer (DiT). Traditional diffusion models typically rely on U-Net as denoising backbones. Recent works propose Diffusion Transformer (DiT) [23] architecture, which supports variable input resolution and sequence length. The DiT combines the generative power of diffusion models with the flexibility of Transformers [34], enabling better scalability and stronger modeling capacity.
@@
3D-Consistent Pose Representation. Recent methods [15, 30, 38, 17, 4] that employ 2D keypoint-based skeletons as motion controls perform well on simple actions. However, they often struggle under complex studio-level production scenarios, leading to abnormal human kinematics and violation of motion physics. 2D keypoints cannot encode 3D occlusion, making it difficult for models to generate realistic motion videos from incomplete or ambiguous pose signals. Another challenge arises from the discrepancy between the reference image and driving pose, which often differ in body shape or identity. Previous attempts [30, 38] mitigate the discrepancy through 2D skeleton scaling in training or heuristic retargeting in testing, yet 2D-derived adaptation inherently suffers from deformation due to inconsistency with 3D motion dynamics and further amplifies estimation noise and inaccuracies.
We argue that an ideal driving pose representation should perceive depth and occlusion, remain identity-agnostic, and retain the potential to scale to multi-person settings. To achieve that, we employ NLFPose [26] to estimate 3D body keypoints $\mathcal{J}=\{\text{J}_{i,j}\}_{i=1,j=1}^{T,N}$, where $\tilde{\text{J}}_{i,j}$ denotes the predicted 3D joint $j$ at frame $i$, and connect them according to the skeletal topology in 3D space. We discard SMPL mesh-like representations, which are inherently human-specific and offer limited controllability for augmentation or retargeting, often resulting in identity leakage. To retain occlusion and depth cues lost in 2D projection, we render 3D skeletons as cylindrical segments to provide motion-driving signals.
@@
Full-Context Driving Pose Injection. Our model builds upon a DiT-based Image-to-Video (I2V) model [36] and inject pose control signals for motion guidance. The common practice for pose information injection is through channel concat [51, 4] or using a pose adapter [38], where the driving pose sequence is embedded and added to the noisy video latents. We first implement the channel concat approach as illustrated in Figure 4, following [51]. While this method shows decent pose-following capability in generation, we find it tend to generate unnatural human pose when the motion is complex despite conditioned on accurate and unambiguous pose. A typical example is the turning motion, where the model often fails to correctly distinguish between front and back views, generating awkward postures. We assume that this limitation stems from the per-frame addition scheme, which fails to provide sufficient temporal context for motion understanding. Since motion is inherently sequential, many actions can only be interpreted correctly within a temporal context. Therefore, we design a new strategy termed full-context pose injection, which enables the model to reason over the entire pose sequence. Specifically, we directly concatenates conditional pose tokens with noisy video tokens to facilitate spatial-temporal interactions between the two modality. As shown in Figure 4, this strategy handles motion correctly in complex scenarios.
@@
where $\text{shift}_{W}$ is a constant shift magnitude and $W_{\max}$ is the width of the reference image. As show in Figure 3, this design helps the model to distinguish driving pose tokens from reference image tokens and noisy video tokens in conjunction with the modified I2V Mask and reference image token injection strategy, enhancing the overall performance. Considering the modality gap between the noisy video tokens and the driving pose tokens, we observe that the model performs best particularly when the shift distance is a relatively large constant. To accommodate different downsampling ratios for pose context, mean-pooling on the 3D-RoPE [29] frequencies is performed according to the applied ratio.
@@ 3.3 Data Collection
Filtering and Pose Extraction Pipeline. We develop a curated pipeline to construct a high-quality dataset for character animation. To meet the requirements for generating studio-grade character animation, we have established the following criteria for data filtering: (1) the character should be the primary focus of each frame, (2) the person should exhibit explicit motion, and (3) the pose should be visually complete, covering either the upper body or the full body. We first detect human presence using YOLO [33] and discard clips where characters are not the main subjects. Next, 2D keypoints are extracted using DWPose [46], retaining only half-body or full-body videos. After this stage, multi-person videos can be further separated based on the number of bounding boxes detected by YOLO [33]. We further employ a large language model [7] to analyze captions, complementing the complex rule-based logic of multi-person motion bounding box checking. For multi-person videos, we employ Samurai [45] to track and split the mask of each character, generating multiple single-human video splits. Subsequently, 3D keypoints are extracted using NLFPose [26] for each video that contain one main character. To meet the motion criteria, we compute the motion speed from the estimated 3D keypoints and discard samples with limited motion, resulting in a clean and motion-rich training set. The motion speed $v$ is calculated as:
@@
Finally, The 3D skeletons are represented as cylindrical segments in 3D space and rendered using the Taichi [16] renderer to preserve spatial relationships. For multi-person data, each split is independently processed by NLFPose to estimate 3D skeletons, which are then composed together in 3D space and rasterized on the 2D canvas. This procedure preserves inter-person occlusion relationships, benefiting from NLFPose’s accurate depth estimation. In practice, we find that our proposed multi-stage pipeline provides more accurate estimation than direct multi-person pose estimation methods like PromptHMR [39], especially in wild cases involving complex interaction.
@@
UniAnimate-DiT [38] 17.79 0.637 0.242 362.27 2.5% 1.7% 0.8% 1.7%
VACE [17] 16.73 0.588 0.263 264.71 9.2% 14.2% 18.3% 32.5%
Wan-Animate [4] 18.54 0.648 0.221 187.61 35.0% 28.3% 24.2% 20.0%
@@
Table 1: Quantitative comparison for SCAIL-14B and baselines. All compared methods are built upon 14B Wan [36] foundation models.
@@ 3.4 Studio-Bench for Evaluation
To evaluate the model capabilities under studio-grade standards, we propose Studio-Bench, a new evaluation benchmark tailored for Studio-Grade character animation. Previous evaluations [44, 15, 52, 49, 31, 38] primarily focus on simple actions, which fail to capture the challenges present in film production, such as complex and dynamic motions, multi-person interactions, and cross-domain animations. Specifically, our benchmark consists of two parts: Self-Driven and Cross-Driven. The first part evaluates motion adherence under complex actions. We examine whether the generated videos maintain correct body structure during large-scale motions, and whether the model can accurately capture inter-person relationships in multi-character scenes, totally containing 130 clips. Quantitative metrics can be computed by directly comparing generated results with paired ground-truth videos for this subset. The second part measures the model’s transferability when the reference image and driving video differ in identity or domain. We manually select a set of real reference images to test generalization, and further construct additional reference images using an image generation tool [27] to introduce variations in style and body shape, obtaining 120 single-character pairs and 10 multi-character pairs. As we simulate in-the-wild reference-driving discrepancy, pose retarget should be enabled for single-character pairs in this subset. We incorporate the retarget logic of Unianimate-dit [38] for VACE [17] which do not natively support retargeting. For this subset, we conduct user study to assess the generation quality. All quantitative results on the subset are based on the 120 single-character pairs to ensure fairness considering some baselines [35] are incompatible with multi-character settings. Evaluation samples are strictly excluded from the training data.